# Coherent Optical Control of Electron Dynamics in Patterned Graphene Nanoribbons

### v1

@@ I Introduction
Coherent electronics is an emerging field which aims to transfer quantum optical functionalities to the solid state domain, see e.g. Higuchi et al. (2017); Heide et al. (2021); Kira and Huber (2025). In contrast to this goal, standard electronics is based on voltage control of electron and hole currents Hori (1997); Cao et al. (2023). In this limit of electron dynamics, properties such as inter-band coherence, are negligible Ashcroft and Mermin (1976); Kohn and Luttinger (1957). However, coherent electronics utilizes quantum coherence as a central ingredient, such as superposition and entanglement, that can be used for a wide range of quantum technological applications Dowling and Milburn (2003).
Light-driven graphene nanoribbons are promising candidates for coherent electronics Higuchi et al. (2017); Heide et al. (2021); Ojeda Collado et al. (2025), due to their large electro-optical response, high electron mobility, and the general robustness of graphene Castro Neto et al. (2009); García de Abajo (2014); Bonaccorso et al. (2010). In particular, armchair graphene nanoribbons exhibit an energy band gap in their electronic band structure that can be experimentally tuned by adjusting the width of the nanoribbon Castro Neto et al. (2009); Zheng et al. (2007). These nanoribbons can be synthesized with atomic precision Verzhbitskiy et al. (2016); Wang et al. (2021) and used to design electronic devices Ojeda Collado et al. (2025). Control of electron dynamics in graphene nanoribbons has been reported in Refs. Jensen et al. (2013); Perez-Piskunow et al. (2014); Babajanov et al. (2014); Higuchi et al. (2017); Koppens et al. (2014); Wang et al. (2021); Singh et al. (2025).
@@
To illustrate our methodology, we show the electron dynamics for optimal $\pi/2$ and $\pi$ pulses and demonstrate that a second pulse can be used to reset the system to its initial state, recovering the equilibrium electronic distribution in a coherent manner. Finally, given the tremendous progress in time and angle-resolved photoemission spectroscopy (trARPES) Freericks et al. (2009); Sprinkle et al. (2009); Boschini et al. (2024); Merboldt et al. (2025), we simulate trARPES of the system under the action of these pulses and conclude that this technique can be used to directly test our predictions experimentally.
This paper is organized as follows. In Sec. II we discuss how to engineer the band structure of the patterned nanoribbon to support quantum optical control. In Sec. III we include the coupling to the electromagnetic field, which we optimize in Sec. IV to engineer a Rabi Hamiltonian. We demonstrate in Sec. V coherent control of the electron dynamics of graphene nanoribbons. In particular, we show that a pi pulse inverts the band population of the valence band and conduction band, and a pi/2 pulse prepares a superposition state where an alternating photocurrent is generated. Finally, in Sec. VI, we determine the electron distribution $n(k,\omega)$ which can be measured in time-resolved ARPES experiments.
@@ II Band structure engineering using gate voltages
We illustrate how the gate voltage pattern generates the desired flat band structure in Fig. 2, where we show the band structure associated with the Hamiltonian Eq. (3) for zero and high gate voltage, U=100VU=100\ $\mathrm${V}\$. As the gate voltage $U$ increases, the lowest-energy bands, plotted in red, flatten and the band gap decreases slightly. For sufficiently high gate voltages, the band gap converges to $\Delta=1.104\ $\mathrm{e}\mathrm{V}$\$, see Appendix A. To quantify the effect of the gate voltage on the flattening of the band structure, we define the band flatness $f$ of the $m$th band with energy $\varepsilon_{m}(k)$ as the ratio between the minimal band gap $\Delta_{m}$ and the bandwidth $W_{m}$ Lee et al. (2016),
@@ III Coupling to the electromagnetic field
We now include the coupling to the electromagnetic field via the Peierls substitution Peierls (1933), diagonalize the $42\times 42$ Hamiltonian and reduce it to an effective $2\times 2$ Hamiltonian by considering only the two lowest energy bands which can be flattened by increasing the gate voltage $U$.
@@
where $\phi_{ij}(\boldsymbol{A})=-\frac{e}{\hbar}\int_{\boldsymbol{r}_{j}}^{\boldsymbol{r}_{i}}\boldsymbol{A}d\boldsymbol{r}$ Peierls (1933), is the Peierls phase acquired by the electrons when tunneling from lattice site $j$ to lattice site $i$ in the presence of a vector potential $\boldsymbol{A}=(A_{x},A_{y})$.
@@ VI Time-resolved ARPES
In this section, we determine the signature of the electron distribution $n(k,\omega)$ in a time-resolved ARPES (trARPES) experiment, for the two $\pi$ pulse protocol. This corresponds to measurement of the electron distribution that is Fourier broadened due to the finite pulse length. We calculate the trARPES signature of the electron distribution as Freericks et al. (2009); Broers and Mathey (2022); Nuske et al. (2020); Broers and Mathey (2021)
@@ VII Conclusion
While we work with the specific nanoribbon design, shown in Fig. 1, we emphasize that a similar approach can be performed for a wide range of nanoribbon sizes, and electrode configurations. Moreover, an approach without the proposed electrostatic potential pattern is also feasible, by using graphene nanoribbons with rhomboidal pattern and superzigzag edges, which can be synthesized with atomic precision Beyer et al. (2019). Different variants of graphene nanoribbons with different patterns can also be synthesized Chen et al. (2020); Wang et al. (2021); Alcón et al. (2025), to which our driving field optimization methodology can also be applied.

### v2

@@ I Introduction
Coherent electronics is an emerging field which aims to transfer quantum optical functionalities to the solid state domain, see e.g. [1, 2, 3]. In contrast to this goal, standard electronics is based on voltage control of electron and hole currents [4, 5]. In this limit of electron dynamics, properties such as inter-band coherence, are negligible [6, 7]. However, coherent electronics utilizes quantum coherence as a central ingredient, such as superposition and entanglement, that can be used for a wide range of quantum technological applications [8].
Light-driven graphene nanoribbons are promising candidates for coherent electronics [1, 2, 9], due to their large electro-optical response, high electron mobility, and the general robustness of graphene [10, 11, 12]. In particular, armchair graphene nanoribbons exhibit an energy band gap in their electronic band structure that can be experimentally tuned by adjusting the width of the nanoribbon [10, 13]. These nanoribbons can be synthesized with atomic precision [14, 15] and used to design electronic devices [9]. Control of electron dynamics in graphene nanoribbons has been reported in Refs. [16, 17, 18, 1, 19, 15, 20].
@@
To illustrate our methodology, we show the electron dynamics for optimal $\pi/2$ and $\pi$ pulses and demonstrate that a second pulse can be used to reset the system to its initial state, recovering the equilibrium electronic distribution in a coherent manner. Finally, given the tremendous progress in time and angle-resolved photoemission spectroscopy (trARPES) [21, 22, 23, 24], we simulate trARPES of the system under the action of these pulses and conclude that this technique can be used to directly test our predictions experimentally.
This paper is organized as follows. In Sec. II we discuss how to engineer the band structure of the patterned nanoribbon to support quantum optical control. In Sec. III we include the coupling to the electromagnetic field, which we optimize in Sec. IV to engineer a Rabi Hamiltonian. We demonstrate in Sec. V coherent control of the electron dynamics of graphene nanoribbons. In particular, we show that a $\pi$ pulse inverts the band population of the valence band and conduction band, and a $\pi/2$ pulse prepares a superposition state where an alternating photocurrent is generated. Finally, in Sec. VI, we determine the electron distribution $n(k,\omega)$ which can be measured in time-resolved ARPES experiments.
@@ II Band structure engineering using gate voltages
We illustrate how the gate voltage pattern generates the desired flat band structure in Fig. 2, where we show the band structure associated with the Hamiltonian Eq. (3) for zero and high gate voltage, U=100VU=100\ $\mathrm${V}\$. As the gate voltage $U$ increases, the lowest-energy bands, plotted in red, flatten and the band gap decreases slightly. For sufficiently high gate voltages, the band gap converges to $\Delta=1.104\ $\mathrm{e}\mathrm{V}$\$, see Appendix A. To quantify the effect of the gate voltage on the flattening of the band structure, we define the band flatness $f$ of the $m$th band with energy $\varepsilon_{m}(k)$ as the ratio between the minimal band gap $\Delta_{m}$ and the bandwidth $W_{m}$ [25],
@@ III Coupling to the electromagnetic field
We now include the coupling to the electromagnetic field via the Peierls substitution [26], diagonalize the $42\times 42$ Hamiltonian and reduce it to an effective $2\times 2$ Hamiltonian by considering only the two lowest energy bands which can be flattened by increasing the gate voltage $U$.
@@
where $\phi_{ij}(\boldsymbol{A})=-\frac{e}{\hbar}\int_{\boldsymbol{r}_{j}}^{\boldsymbol{r}_{i}}\boldsymbol{A}d\boldsymbol{r}$ [26], is the Peierls phase acquired by the electrons when tunneling from lattice site $j$ to lattice site $i$ in the presence of a vector potential $\boldsymbol{A}=(A_{x},A_{y})$.
@@ VI Time-resolved ARPES
In this section, we determine the signature of the electron distribution $n(k,\omega)$ in a time-resolved ARPES (trARPES) experiment, for the two $\pi$ pulse protocol. This corresponds to measurement of the electron distribution that is Fourier broadened due to the finite pulse length. We calculate the trARPES signature of the electron distribution as [21, 27, 28, 29]
@@ VII Conclusion
While we work with the specific nanoribbon design, shown in Fig. 1, we emphasize that a similar approach can be performed for a wide range of nanoribbon sizes, and electrode configurations. Moreover, an approach without the proposed electrostatic potential pattern is also feasible, by using graphene nanoribbons with rhomboidal pattern and superzigzag edges, which can be synthesized with atomic precision [30]. Different variants of graphene nanoribbons with different patterns can also be synthesized [31, 15, 32], to which our driving field optimization methodology can also be applied.